OpenStudy (anonymous):

Mary purchased a package of 24 different plants, but she only needed 20 plants for planting. In how many ways can she select the 20 plants from the package to be planted? please show work!

OpenStudy (dumbcow):

use combinations \[\left(\begin{matrix}24 \\ 20\end{matrix}\right) = \frac{24!}{4!20!}\]

OpenStudy (anonymous):

it is \[\frac{24\times 23\times 22\times 21}{4\times 3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

in particular since \[4\times 3\times 2=24\] you get \[23\times 22\times 21\]

OpenStudy (nilankshi):

its rite

OpenStudy (anonymous):

we then put \[3^x=\frac{3+\sqrt{57}}{2}\]

OpenStudy (anonymous):

you can ignore the negative answer because hold on what is your original equation?

OpenStudy (anonymous):

i think order matters hence its a permutations problem 24p20